At the beginning of 2019, Metatec Inc. acquired Ellison Technology Corporation for $560 million. In addition to cash, receivables, and inventory, the following assets and their fair values were also acquired:

Plant and equipment (depreciable assets) $ 146 million
Patent 36 million
Goodwill 120 million

The plant and equipment are depreciated over a 10-year useful life on a straight-line basis. There is no estimated residual value. The patent is estimated to have a 5-year useful life, no residual value, and is amortized using the straight-line method.

At the end of 2021, a change in business climate indicated to management that the assets of Ellison might be impaired. The following amounts have been determined:

Plant and equipment:
Undiscounted sum of future cash flows $ 76 million
Fair value 56 million
Patent:
Undiscounted sum of future cash flows $ 20 million
Fair value 13 million
Goodwill:
Fair value of Ellison Technology Corporation $ 406 million
Fair value of Ellison's net assets (excluding goodwill) 350 million
Book value of Ellison's net assets (including goodwill) 430 million*

*After first recording any impairment losses on plant and equipment and the patent.

4. Determine the amount of any impairment loss to be recorded, if any, for the three assets.

Determine the amount of any impairment loss to be recorded, if any, for the three assets. (Enter your answers in millions. Negative amounts should be indicated by a minus sign.)

Impairment Loss
Plant and equipment million
Patent million
Goodwill million
Answer #1

wwImpairment loss will be reconized when sum of future cash flows is less than book value

Plant

Book Value at the end of year 2021:

Dep annual as per staight line method= 146/10= 14.6m

for 3 years = 14.6*3= 43.8m

(2019,2020and 2021)

book value of plant as end of 2021= 146-43.8= 102.2m

Impairment loss will be reconized for plant as the sum of future cash flows is less than book value

therefore impairment loss= Book value- fair value= 102.2-56= $46.2 m

Patent

Book Value at the end of year 2021:

Dep annual as per staight line method= 36/5= 7.2

for 3 years= 7.2*3= 21.6

book value of patent as end of 2021= 36-21.6= $14.4m

Impairment loss will NOT be reconized for patent as the sum of future cash flows ($20) is more than book value ($14.4)

Goodwill

implied fair value of goodwill = 406-350= 56m

carrying value of goodwill = 120m

goodwill impairemnt= 120-56= 64m